ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 51, 52 Crespo Crespo, Cecilia; Pizzo, Aldo; Ponteville, Christiane; Villella, José - Universidad Nacional de General San Martín, Buenos Aires, Argentina Thinking about Geometry Teaching In the scope of a curricular transformation in teaching in Argentina, it appears as a necessity to revise the existing teaching practice in the area of Geometry, as oriented towards students from 12 to 17 years old. Such reflexion ground for teachers about certain concepts like tridimensional space and plane should involve both a disciplinar and didactic up-dating. It might be carried away thought lectures and it requires of a working module such as the one that represents the backbone of this presentation. This work is structured upon the resolution, analysis, discussion and writing activities that, centered around disciplinary contents it permits the reader to meditate about the didactic transposition of the contents hereby reelaborated. Reference: J. Villella, C. Crespo Crespo, Ch. Ponteville: Cuando la Geometría es el tema de la reflexión matemática. Universidad Nacional de General San Martín, Buenos Aires, Argentina. (1998)